# Localizing Before Answering: A Hallucination Evaluation Benchmark for Grounded Medical Multimodal LLMs

### v3

@@ 1 Introduction
Large Multimodal Models (LMMs), or Multi-modal Large Language Models (M-LLMs) Alayrac et al. (2022); Bai et al. (2023); Liu et al. (2024b) such as GPT-4V OpenAI et al. (2023) achieve superb performance in understanding data from multiple modalities, such as vision, and language, and generating human-like texts. Leveraging the strong capabilities of those foundational models, recent works Li et al. (2024); Chen et al. (2024); Zhang et al. (2023a); Chen et al. (2023); Moor et al. (2023); Wu et al. (2023b, a); Yan and Pei (2022) have developed M-LLMs for medical imaging applications such as medical Visual Question Answering (VQA). Recently, many multimodal foundation models (LLaVA-Med Li et al. (2024), CheXagent Chen et al. (2024), GPT-4 OpenAI et al. (2023), Gemini Team et al. (2023)) have emerged and demonstrated impressive reasoning and comprehension capabilities on biomedical queries and related domains. These models can assist clinical professionals in interpreting and gaining insights on medical images, helping the diagnosis and treatment processes be more efficient.
@@
Figure 1: An example of Med-VQA from our dataset. We follow LVLM-Interpret Stan et al. (2024) to visualize a) the attention map of the image with respect to the model response, and b) the relevancy score between the answer and input words. LLaVA-Med Li et al. (2024), as reference model fails to attend to the region of interest, and its answer has a low relevancy score to the image. Several hallucination reduction methods can generate responses more relevant to the image, but may not localize the target region.
@@
Due to the sensitive nature of the biomedical domain, the accuracy and trustworthiness of these foundational models are important. Though impressive, Large Language Models (LLMs) and multimodal LLMs are prone to hallucinate contents. Specifically, the models produce false answers, which are not grounded in the visual evidence. Given a question such as “Does the patient have pneumonia in his right lung?”, many LLMs are prone to answer “Yes” without performing a deep analysis of the image content. This behavior arises because the observation pneumonia frequently co-occurs with parts of the lung, leading to shallow correlations rather than rigorous reasoning. Fig. 1 illustrates the answer of some open-source LMMs, as well as the visual cross-attention weight of the models during generation. These models tend to hallucinate and attend to irrelevant image regions when answering these questions.
@@
This study shows two types of shortcut learning that current multimodal LLMs suffer from: textual and visual shortcut learning. First, the models attend to non-important text tokens with higher relevancy score than any image tokens. As shown in Fig. 1b, the model produces higher relevancy scores on token lung than any image token. Second, on the image, the model attends to non-queried image regions (as shown in Fig. 1a). This shows the model learns shortcut on irrelevant text and image tokens, instead of basing on the visual evidence when answering the questions.
@@
To address these concerns, we introduce a new benchmark, dubbed HEAL-MedVQA (Hallucination EvAluation via Localization in Medical VQA) to dissect the LLM’s ability to localize at visual evidence when answering. Specifically, this paper proposes two evaluation protocols, called Textual Perturbation Test (TPT), and Visual Perturbation Test (VPT), which probe LLMs’ sensitivity towards textual and visual shortcuts. To support this evaluation, we curate adversarial VQAs and procure pixel-level annotations of affected anatomy regions from three board-certified radiologists from two large-scale public datasets, MIMIC-CXR Johnson et al. (2019) and VinDr-CXR Nguyen et al. (2020). In total, we construct over 60,000 question-answer pairs probing LLM’s textual and visual shortcut learning. Our curated medical benchmark provides structured codebase, facilitating comprehensive comparisons of advanced LLMs. The code and dataset will be released at https://loba-ijcai25.github.io.
@@
To this end, we propose a Localize-before-Answer (LobA) framework that forces LMM to generate visual-evidence tokens, which is then input to segmentation decoder to localize the area of interests before answering. Based on the localization map, LobA self-prompts to enhance the model’s attention, forcing it to leverage visual evidence relevant to the question before answering. In summary, our main contributions are as follows.
@@
  • We introduce HEAL-MedVQA (Hallucination Evaluation via Localization in Medical VQA) benchmark consists of 67K QA pairs, and doctor-annotated segmentation masks with two new evaluation protocols, assessing LLMs’ robustness to textual and visual shortcut learning. Our benchmark includes a well-structured codebase, and comprehensive comparisons of 8 state-of-the-art LLMs and hallucination techniques, thus standardizing hallucination and accuracy evaluations of future multi-modal LLMs.
@@
  • We propose the Localize-before-Answer (LobA) framework, which forces the model to localize the affected anatomical regions and self-prompt to enhance the attention on the target regions before answering.
@@ 2 Related works
Medical Visual Question Answering. Several benchmarks in Med-VQA propose specialized datasets Hasan et al. (2018); Abacha et al. (2019, 2020, 2021). VQA-RAD Lau et al. (2018) offers both open-ended and closed-ended questions. SLAKE Liu et al. (2021) develops bilingual dataset with knowledge graph. Overall, existing benchmarks use accuracy as the main metrics, which overlooks the hallucination evaluation of multi-modal LLMs.
A majority of Med-VQA models are finetuned from other VLMs, namely LLaVA-Med Li et al. (2024) from LLaVA Liu et al. (2024b), CheXagent Chen et al. (2024) from BLIP-2 Li et al. (2023a), or Zhang et al. (2023a); Chen et al. (2023); Moor et al. (2023); Wu et al. (2023b, a); Yan and Pei (2022). Recent approaches, which localize relevant regions in Med-VQA have gained attention in recent times Tascon-Morales et al. (2023, 2024); Zhang et al. (2024a). However, during inference time, these VQA models require human-annotated masks to localize region of interests. In contrast, our method explicitly learns to localize during training, bypassing the needs to acquire human annotations when inferring.
Hallucination in VQA. To overcome hallucination, there are two main approaches. First, data-centric methods focus on augmenting existing VQA training datasets and propose new benchmarks for evaluation. In Li et al. (2023b) and Liu et al. (2023), they show that frequently appeared objects tend to guide models’ answers during inference time more than absent or less frequent ones. As such, POPE Li et al. (2023b) adds adversarial questions asking for the presence of objects that are absent on the image. LRV-Instruction Liu et al. (2023) introduces negative instructions which are irrelevant to the visual content.
Second, model-centric methods focus on modifying the model architecture to support the visual component. Intern-VL Chen and others (2024), and LLaVA-1.5-HD Liu et al. (2024a) scale up the vision foundation model to allow better visual understanding. Others, such as He et al. (2024); Jain et al. (2024) append several levels of information (segmentation mask, object list, task-based knowledge) to enrich the visual context. Lastly, several methods proposes inference-based methods, which augments the inference of the pre-trained LLMs. VCD Leng et al. (2024), and M3ID Favero et al. (2024) aim to ground inference by negating a contrastive version of the input where the image is distorted or dropped. OPERA Huang et al. (2024) discourages overly confident but erroneous information accumulated in self-attention layers by applying a penalty. Overall, existing hallucination techniques are applied on the image level, while neglecting region-level details. In contrast, we propose to explicitly reason on the pixel level before answering.
@@ 3.1 New Evaluation Protocol — Sensitivity Test
Most current VQA works Abacha et al. (2019); Liu et al. (2021); Lau et al. (2018) primarily rely on the accuracy of the model’s output answers as the key evaluation metric. However, these metrics overlook other bias factors that could influence the model’s responses. In particular, we aim to assess two key phenomena in medical LMMs:
@@
  • Language Shortcut: Despite generating correct answers, the model relies on language shortcut learnt in the training dataset, rather than the actual visual content. For example, given a question “Does the Right lower lung have lung opacity”, models can answer “Yes” due to the co-occurence of the words “lung” and “lung opacity” in the training dataset.
@@
  • Vision Shortcut: The model might concentrate on non-relevant regions (as in Fig. 1a). This behavior results in the unreliability of LMMs, even when providing correct answers, as they fail to rely on medical visual evidence.
@@
To that end, we propose two new evaluation protocols called Textual Perturbation Test (TPT), and Visual Perturbation Test (VPT), as shown in Fig. 2.
Figure 2: Textual and Visual Perturbation Tests for our experiments. For Textual Perturbation Test (TPT), we randomly replace either the anatomy or the disease in the original question with another so that the new anatomy/disease pair is logical (present in the training dataset). For Visual Perturbation Test (VPT), we replace the visual area of interest (e.g. left lower lung) with the segmented area of the same region from another image that does not share the same diagnosis as the current one. We measure the percentage of instances a model changes the response from “Yes” to “No” to evaluate the model’s sensitivity to visual and language shortcuts.
@@
Textual Perturbation Test. In disease diagnosis, medical LLMs can exploit the co-occurrence of disease and commonly affected regions to provide superficial answers. To assess how sensitive a model is to these language biases, we propose to swap the key entities, i.e., anatomy or disease, and evaluate the model’s sensitivity to the perturbed questions. The evaluation protocol is as follows.
@@
  • From a test set of Yes/No binary questions under the form of “Does anatomy have disease”, we collect the True Positive samples of the model (samples correctly predicted as “Yes”).
@@
  • We randomly replace either the anatomy or disease with another new anatomy/or disease. The new anatomy and disease pair has to be logical, which means they does co-occur in the dataset. For example, in Fig. 2 (Above), the question “Does left lower lung suffer from pneumonia?” can be replaced by either switching the anatomy to right upper lung, or the disease to pulmonary fibrosis.
  • We measure the percentage of the answers being changed from “Yes” to “No” when applying this protocol as the textual perturbation test (TPT) score.
@@
Visual Perturbation Test. We propose to assess whether the model grounds the answer on the anatomical region of interests. Specifically, when querying the presence of a disease in a particular area, e.g., the left lower lung, we take a segmented region of the left lower lung from another image that does not share the same diagnosis as the current one. This segmented region is resized and blended onto the original image. By doing so, we expect the model to focus its attention on the replaced region of interest to produce the correct diagnosis, altering the response from “Yes” to “No.” We define the visual perturbation test (VPT) score as the percentage of answers that change from “Yes” to “No” under this protocol.
@@ 3.2 Dataset Curation
To assess the LLM’s robustness against both textual and visual shortcuts, we introduce a new benchmark called HEAL-MedVQA (Hallucination Evaluation via Localization in Medical VQA). Built upon the two large scale radiology datasets, VinDr-CXR Nguyen et al. (2020) and MIMIC-CXR Johnson et al. (2019), we construct HEAL-MedVQA by mapping spatial relationships between diseases and anatomical locations, and probing the model’s robustness towards disease and location perturbation on both texts and images. Fig. 4 illustrates our overall pipeline of data construction. We extract the disease bounding boxes and anatomical masks from the image, and based on the overlap between disease and anatomy localization, we determine their spatial relationships. Then several QA templates are used to generate VQA samples.
Figure 3: Data processing pipeline. Our data processing pipeline consists of 4 main steps: 1. Anatomy Segmentation., where we obtain segmentation masks of core anatomies; 2. Disease Extraction., where the disease labels and their bounding boxes are obtained through either a Detection model, or from radiologists’ annotation; 3. Anatomy-Disease Relation Mapping, where anatomies are mapped with diseases based on their IoU score; 4. QA Generation, where question-answer pairs are generated based on the anatomy-disease mappings.
For the disease bounding boxes, VinDr-CXR database provides bounding box annotations for 22 medical findings. We then train a standard detection model, YOLOv5 Jocher et al. (2020), on VinDR-CXR dataset and obtain the bounding boxes on MIMIC-CXR. To reduce the false positive detection, we extract the disease labels from the associated medical reports, and filter out diseases which are absent on that patient.
For anatomy segmentation, we define eight main anatomies, e.g. Right lower lung, Left middle lung, Aorta, Heart, Trachea, where chest diseases commonly occur. On the training dataset, we then obtain the segmentation mask by using a model pre-trained on PAXRay++ dataset Seibold and others (2022). On the test datasets, we consult three board-certified radiologists to annotate 8 main anatomical structures on the pixel level. By providing doctor-annotated pixel-level segmentation, this research offers a new benchmark to evaluate the visual reasoning and the robustness to hallucination of current Multimodal LLMs.
@@
After obtaining a list of disease-anatomy mappings, for each image we randomly generate 2-5 questions and answers pairs based on these mappings. We focus on two main question formats: Close-ended questions and Open-ended questions, as exemplified by Fig. 4 Close-ended questions are binary questions of formats such as “Does anatomy have disease”, “Is there disease present in anatomy”. Based on the disease-anatomy mappings, the close-ended questions is either of type Positive, which queries the disease that is present in the anatomy, or of type Hallucinated, which queries the disease that is not present in the anatomy.
Open-ended questions require the model to makes more comprehensive diagnosis. It includes two types of questions: normal and abnormality-query questions which respectively probe regions without and with diseases. Sample question templates include “Are there any abnormalities in the heart?” or “List all abnormalities in the Aorta”.
@@
In total, our HEAL-MedVQA dataset comprises 45,331 closed-ended questions and 22,573 open-ended ones. We compare our benchmark with those previously reported at Table 1. To the best of our knowledge, we are the first benchmark providing large-scale dataset with comprehensive hallucination metrics, and doctor-annotated masks to evaluate the hallucination of emerging LLMs.
Datasets # Images # QA pairs

### v4

# Localizing Before Answering: A Benchmark for Grounded Medical Visual Question Answering

@@ 1 Introduction
Large Multimodal Models (LMMs), or Multi-modal Large Language Models (M-LLMs) Alayrac et al. (2022); Bai et al. (2023); Liu et al. (2024b) such as GPT-4V OpenAI et al. (2023) achieve superb performance in understanding data from multiple modalities, such as vision and language, and generating human-like texts. Leveraging the strong capabilities of those foundational models, recent works Li et al. (2024); Chen et al. (2024); Zhang et al. (2023a); Chen et al. (2023); Moor et al. (2023); Wu et al. (2023b, a); Yan and Pei (2022) have developed M-LLMs for medical imaging applications such as medical Visual Question Answering (VQA). Recently, many multimodal foundation models (LLaVA-Med Li et al. (2024), CheXagent Chen et al. (2024), GPT-4 OpenAI et al. (2023), Gemini Team et al. (2023)) have emerged and demonstrated impressive reasoning and comprehension capabilities on biomedical queries and related domains. These models can assist clinical professionals in interpreting and gaining insights on medical images, helping the diagnosis and treatment processes be more efficient.
@@
Figure 1: An example of Med-VQA from our dataset. We follow LVLM-Interpret Stan et al. (2024) to visualize a) the attention map of the image with respect to the model response, and b) the relevancy score between the answer and input words. LLaVA-Med Li et al. (2024), as reference model, fails to attend to the region of interest, and its answer has a low relevancy score to the image. Several hallucination reduction methods can generate responses more relevant to the image, but may not localize the target region.
@@
Due to the sensitive nature of the biomedical domain, the accuracy and trustworthiness of these foundational models are important. Though impressive, Large Language Models (LLMs) and multimodal LLMs are prone to hallucinating contents. Specifically, the models produce false answers, which are not grounded in the visual evidence. Given a question. “Does the patient have pneumonia in his right lung?”, many LLMs are prone to answer “Yes” without reasoning the image content. This behavior arises because the observation of pneumonia frequently co-occurs with parts of the lung, leading to spurious correlations instead of rigorous reasoning. Fig. 1 illustrates the answer of some open-source LMMs, and the cross-attention map of the models during generation. These models tend to hallucinate and attend to irrelevant image regions when answering these questions.
@@
This study shows two types of shortcut learning that current multimodal LLMs suffer from: textual and visual shortcut learning. First, the models attend to non-important text tokens with higher relevancy score than any image tokens. As shown in Fig. 1b, the model produces higher relevancy scores on the token lung than on any image token. Second, on the image, the model attends to non-queried image regions (as shown in Fig. 1a). This shows the model learns shortcuts on irrelevant text and image tokens, instead of basing on the visual evidence when answering the questions.
@@
To address these concerns, we introduce a new benchmark, dubbed HEAL-MedVQA (Hallucination EvAluation via Localization in Medical VQA), to dissect the LLM’s ability to localize at visual evidence when answering. Specifically, this paper proposes two evaluation protocols, called Textual Perturbation Test (TPT) and Visual Perturbation Test (VPT), which probe LLMs’ sensitivity towards textual and visual shortcuts. To support this evaluation, we curate adversarial VQAs and procure pixel-level annotations of affected anatomy regions from three board-certified radiologists from two large-scale public datasets, MIMIC-CXR Johnson et al. (2019) and VinDr-CXR Nguyen et al. (2020). In total, we construct over 60,000 question-answer pairs probing LLM’s textual and visual shortcut learning. Our curated medical benchmark provides a structured codebase, facilitating comprehensive comparisons of advanced LLMs.
@@
To this end, we propose Localize-before-Answer (LobA), a framework that forces LMM to generate visual-evidence tokens, which is then input to the segmentation decoder to localize the area of interest before answering. Based on the localization map, LobA self-prompts to enhance the model’s attention, forcing it to leverage visual evidence relevant to the question before answering. In summary, our main contributions are as follows.
@@
  • We introduce HEAL-MedVQA (Hallucination Evaluation via Localization in Medical VQA), a benchmark that consists of 67K QA pairs, and doctor-annotated segmentation masks with two new evaluation protocols, assessing LLMs’ robustness to textual and visual shortcut learning. Our benchmark includes comprehensive comparisons of 8 state-of-the-art LLMs and hallucination techniques, thus standardizing hallucination and accuracy evaluations of future multi-modal LLMs.
@@
  • We propose the Localize-before-Answer (LobA) framework, which forces the model to localize the affected anatomical regions and self-prompt to enhance attention on the target regions before answering.
@@ 2 Related Works
Medical Visual Question Answering. Several benchmarks in Med-VQA propose specialized datasets Hasan et al. (2018); Abacha et al. (2019, 2020, 2021). VQA-RAD Lau et al. (2018) offers both open-ended and closed-ended questions. SLAKE Liu et al. (2021) develops a bilingual dataset with a knowledge graph. Overall, existing benchmarks use accuracy as the main metric, which overlooks the hallucination evaluation of multi-modal LLMs.
A majority of Med-VQA models are finetuned from other VLMs, namely LLaVA-Med Li et al. (2024) from LLaVA Liu et al. (2024b), CheXagent Chen et al. (2024) from BLIP-2 Li et al. (2023a), or Zhang et al. (2023a); Chen et al. (2023); Moor et al. (2023); Wu et al. (2023b, a); Yan and Pei (2022); Huy et al. (2025). Recent approaches, which localize relevant regions in Med-VQA, have gained attention in recent times Tascon-Morales et al. (2023, 2024); Zhang et al. (2024a). However, during inference time, these VQA models require human-annotated masks to locate regions of interest. In contrast, our method explicitly learns to localize during training, bypassing the need to acquire human annotations when inferring.
Hallucination in VQA. To overcome hallucination, there are two main approaches. First, data-centric methods focus on augmenting existing VQA training datasets and propose new benchmarks for evaluation. In Li et al. (2023b) and Liu et al. (2023), they show that frequently appearing objects tend to guide models’ answers during inference time more than absent or less frequent ones. As such, POPE Li et al. (2023b) adds adversarial questions asking for the presence of objects that are absent from the image. LRV-Instruction Liu et al. (2023) introduces negative instructions, which are irrelevant to the visual content.
Second, model-centric methods focus on modifying the model architecture to support the visual component. Intern-VL Chen and others (2024), and LLaVA-1.5-HD Liu et al. (2024a) scale up the vision foundation model to allow better visual understanding. Others, such as He et al. (2024); Jain et al. (2024) append several levels of information (segmentation mask, object list, task-based knowledge) to enrich the visual context. Lastly, several methods propose inference-based methods, which augment the inference of the pre-trained LLMs. VCD Leng et al. (2024) and M3ID Favero et al. (2024) aim to ground inference by negating a contrastive version of the input where the image is distorted or dropped. OPERA Huang et al. (2024) discourages overly confident but erroneous information accumulated in self-attention layers by applying a penalty. Overall, existing hallucination techniques are applied on the image level, while neglecting region-level details. In contrast, we propose to explicitly reason on the pixel level before answering.
@@ 3.1 New Evaluation Protocol — Sensitivity Test
Conventional accuracy metrics Abacha et al. (2019); Liu et al. (2021); Lau et al. (2018) overlook bias factors influencing the model’s responses. This paper assesses two key phenomena in medical LMMs:
@@
  • Language Shortcut: Despite generating correct answers, the model relies on language shortcuts learnt in the training dataset, rather than the actual visual content. For example, given a question “Does the right lower lung have pneumonia?”, models can answer “Yes” due to the co-occurrence of lung and pneumonia in the training set.
@@
  • Vision Shortcut: The model can concentrate on non-relevant regions (as in Fig. 1a). This behavior results in the unreliability of LMMs, even when providing correct answers, as they ignore visual evidence.
@@
To this end, we propose two new evaluation protocols, called Textual Perturbation Test (TPT) and Visual Perturbation Test (VPT), as shown in Fig. 2.
Figure 2: Textual and Visual Perturbation Tests. TPT: Replace the anatomy or disease term in the question. VPT: Swap the region of interest with the same region from an image of a different class. We report the percentage of “Yes” → “No” flips to assess model sensitivity to shortcuts.
@@
Textual Perturbation Test. To assess how sensitive a model is to language biases, we propose to swap the key entities, i.e., anatomy or disease, and evaluate the model’s sensitivity to the perturbed questions. The evaluation protocol is as follows.
@@
  • From the test set of binary questions - “Does anatomy have disease?”, we collect True Positive samples having correct “Yes” predictions.
@@
  • We randomly replace either the anatomy or disease with another valid co-occurring term. For example, in Fig. 2 (Above), the question “Does left lower lung suffer from pneumonia?” is replaced by either switching the anatomy to right upper lung, or to another disease.
  • The percentage of “Yes” → “No” flips defines the Textual Perturbation Test (TPT) score.
@@
Visual Perturbation Test. We propose to assess whether the model grounds the answer on the anatomical region of interest. Specifically, when querying the presence of a disease in a particular area, e.g., the left lower lung, we take a segmented region of the left lower lung from another image that does not share the same diagnosis as the current one. This segmented region is resized and blended onto the original image. By doing so, we expect the model to focus its attention on the replaced region of interest to produce the correct diagnosis, altering the response from “Yes” to “No”. We define the visual perturbation test (VPT) score as the percentage of answers that change from “Yes” to “No” under this protocol.
@@ 3.2 Dataset Curation
To assess the LLM’s robustness against both textual and visual shortcuts, we introduce a new benchmark called HEAL-MedVQA (Hallucination Evaluation via Localization in Medical VQA). Built upon VinDr-CXR Nguyen et al. (2020) and MIMIC-CXR Johnson et al. (2019), we construct HEAL-MedVQA by mapping spatial relationships between diseases and anatomical locations, and probing the model’s robustness towards disease and location perturbation on both texts and images. Fig. 4 illustrates our overall pipeline of data construction. We extract the disease bounding boxes and anatomical masks from the image. Based on the overlap between disease and anatomy localization, we determine their spatial relationships. Then, several QA templates are used to generate VQA samples. This process can be broken down into four main steps:
Figure 3: Our data processing pipeline.
Anatomy segmentation. We define 8 main anatomies, e.g. Right lower lung, Heart, etc., where chest diseases commonly occur. On the training set, we obtain the segmentation mask using a model pre-trained on PAXRay++ dataset Seibold and others (2022). On the test set, we consult three board-certified radiologists to annotate the segmentation. By providing doctor-annotated pixel-level segmentation, this research offers a new benchmark to evaluate visual reasoning and robustness to hallucination of current M-LLMs.
Disease extraction. For the disease bounding boxes, the VinDr-CXR dataset provides bounding box annotations for 22 medical findings. We train a standard detection model, YOLOv5 Jocher et al. (2020), on the dataset and obtain the bounding boxes on MIMIC-CXR. To reduce false positives, we extract the disease labels from the associated medical reports and filter out absent diseases.
@@
Question Generation. After obtaining a list of disease–anatomy mappings, we randomly generate 2–5 QA pairs per image, focusing on two main formats: close-ended and open-ended, as shown in Fig. 4. Close-ended questions are binary: “Does anatomy have disease?”, “Is there disease present in anatomy?”. Each is either a Positive question, which queries a disease present in the anatomy, or a Hallucinated question, which queries a disease that is not present. Open-ended questions have two types: normal and abnormality-query questions, which respectively probe regions without and with diseases.
@@
In total, our HEAL-MedVQA dataset comprises 45,331 closed-ended questions and 22,573 open-ended ones. We compare our benchmark with those previously reported in Table 1. To the best of our knowledge, we are the first benchmark providing a large-scale dataset with comprehensive hallucination metrics and doctor-annotated masks to evaluate the hallucination of emerging LLMs.
Datasets # Images
# QA
pairs